OpenStudy (anonymous):

To clear fractions from the equation 7 -- n - 1 = -11 2 ---- 4 you could multiply each term on both sides by the number 2. True False

OpenStudy (unklerhaukus):

\[\frac{7}{2}n-1=\frac{-11}{4}\] multiplying by 2 \[2\times\left(\frac{7}{2}n-1\right)=2\times{\frac{-11}{4}}\] \[7n-2={\frac{-11}{2}}\] have all the fractions gone?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so its false

OpenStudy (anonymous):

?